# Enhancing Reusability of Learned Skills for Robot Manipulation via Gaze and Bottleneck

### v2

@@ Abstract
Autonomous agents capable of diverse object manipulations should be able to acquire a wide range of manipulation skills with high reusability. Although advances in deep learning have made it increasingly feasible to replicate the dexterity of human teleoperation in robots, generalizing these acquired skills to previously unseen scenarios remains a significant challenge. In this study, we propose a novel algorithm, Gaze-based Bottleneck-aware Robot Manipulation (GazeBot), which enables high reusability of the learned motions even when the object positions and end-effector poses differ from those in the provided demonstrations. By leveraging gaze information and motion bottlenecks—both crucial features for object manipulation—GazeBot achieves high generalization performance compared with state-of-the-art imitation learning methods, without sacrificing its dexterity and reactivity. Furthermore, the training process of GazeBot is entirely data-driven once a demonstration dataset with gaze data is provided. Videos and code are available at https://crumbyrobotics.github.io/gazebot.
@@ I Introduction
Recent advancements utilizing powerful neural networks such as Transformers have made deep imitation learning increasingly capable of reproducing dexterity to a certain extent [29, 5, 14]. However, significant issues persist regarding their generalization capabilities. Although generalization in object manipulation occurs at multiple levels, even the most fundamental aspects, such as changes in object position and the end-effector pose, are known to cause drastic reductions in success rates with variations of just a few centimeters [4]. For instance, ACT [29], a model recognized for its strong dexterous capabilities, has only been validated with objects placed on white tape with an accuracy of approximately 5 cm. Although ACT demonstrated high success rates under these specific conditions in our experiments, it was unable to reach objects placed in unseen positions (Figure 1), highlighting the poor generalization capabilities acquired through this method.
@@
The purpose of this study is to achieve imitation that can accurately perform demonstrated object manipulations under various object positions and initial end-effector poses. Owing to the limited generalization capability of conventional imitation learning, exhaustive demonstration collection is currently required to ensure that the robot behaves correctly under various object positions and end-effector poses [3, 15, 16]. From a short-term perspective such as automating a certain task, it might be sufficient to collect data until the desired capability is achieved. However, for enabling robots to autonomously and adaptively expand human-aligned behaviors in unknown environments, it is essential that actions learned from limited demonstrations can be reused in slightly different situations. To address this need, we propose Gaze-based Bottleneck-aware Robot Manipulation (GazeBot), an object manipulation imitation method that enables the reuse of acquired skills even with object positions and end-effector poses not included in the provided demonstrations. As illustrated in Figure 1, the object positions and initial end-effector poses used in the demonstrations are restricted to a designated region to evaluate the generalization performance. In this setting, GazeBot can accurately perform the task not only under in-distribution (ID) conditions (i.e., within the designated region) but also under out-of-distribution (OOD) conditions (i.e., outside the designated region).
@@
To reuse skills learned within the ID domain in OOD situations, it is necessary to (1) establish an object representation that is robust to changes in object position and (2) develop an action policy architecture capable of accurate control under previously unseen end-effector poses. To achieve this, we draw inspiration from human object manipulation, where gaze on the target object not only provides a visual representation independent of the object’s absolute position but also exhibits strong gaze–hand coordination during end-effector movements such as reaching [10, 8]. Concretely, we first represent the entire field of view using a 3D point cloud and then crop a cubic region around the gaze position—referred to as gaze-centered point cloud—from the entire point cloud. By using this gaze-centered point cloud as input to the action prediction, we realize an object representation that is robust to variations in object position. Next, based on the action predictivity in the gaze-centered point cloud, we perform a data-driven action segmentation of the overall manipulation into (1) a reaching motion to the vicinity of the gaze position and (2) a gaze-centered dexterous action. We define the temporal boundary between these motions as a bottleneck, and predict the bottleneck pose, the end-effector pose at this bottleneck, from the 3D gaze position and gaze-centered point cloud. This approach allows accurate prediction of the bottleneck pose even for unseen object positions, and this prediction is independent of the current end-effector pose. Consequently, GazeBot can handle various object positions and end-effector poses by first executing a rough reaching motion to the bottleneck pose, then performing a gaze-centered dexterous action using only the gaze-centered point cloud. This approach enables the reuse of the learned motion in unseen conditions. Here, reaching the bottleneck is achieved by generating an end-effector trajectory that smoothly connects the current end-effector pose and the bottleneck pose using a first-order Bézier curve, which provides sufficient expressive power for the reaching motion while avoiding unnecessary complexity that could compromise the model’s generalization. Furthermore, GazeBot updates all actions at every step, and a fully parametric method based on a Transformer is used to directly output subsequent end-effector poses for the gaze-centered dexterous actions. This approach ensures that dexterity and reactivity are not sacrificed in the pursuit of generalization.
In summary, our key results and contributions are as follows:
  • We propose a gaze-centered point cloud as a visual representation that is invariant to the object position.
  • We introduce a method for data-driven action segmentation based on bottleneck determination that clarifies which actions can be reused.
  • We implement a reaching method to the bottleneck pose that remains accurate even with unseen object positions and end-effector poses.
  • We develop a novel imitation learning model, GazeBot, which demonstrates superior generalization performance compared with state-of-the-art models by leveraging gaze and bottlenecks.
@@ II Related Work
Gaze-based Object Manipulation. Some works have previously proposed imitation learning methods using gaze data collected from a remote human operator during teleoperated demonstrations for action prediction [12, 13, 14]. These gaze-based methods have previously exhibited advantages such as enhanced robustness by disregarding task-irrelevant objects [12] and improved dexterity by focusing on task-relevant regions of visual inputs [13, 14]. However, because these methods rely on image cropping for gaze-centered images, they are susceptible to visual variations caused by changes in object position. In this study, we employ a gaze-centered point cloud that is robust to positional changes, and our method further improves the reusability of acquired skills for unseen object positions and end-effector poses, in addition to the conventional benefits of gaze.
@@
Data-driven Action Segmentation. Segmenting actions into reaching motions and dexterous actions has been proposed to improve dexterity [13], increase success rates for long-horizon tasks [14, 1, 22], and enable high generalization capabilities [11, 22]. Kim et al. proposed data-driven segmentation methods based on end-effector velocity [13] or the visibility of the end-effector within a gaze-centered image [14]. However, these approaches often fail in tasks where high dexterity is not required or when the end-effector is not visible during dexterous actions, such as manipulating with a long stick. In contrast, our approach segments motions at bottlenecks, which are determined based on action predictivity in the gaze-centered point cloud. This action predictivity-based approach offers a more general and data-driven segmentation scheme compared with these previous methods.
@@
Generalization to Out-of-Distribution. Generalization to out-of-distribution scenarios is required across diverse levels and factors [2], including adaptation to a variety of object poses and unseen objects in the same category [20, 24, 7], to changes in the environment such as varying backgrounds, camera positions, or distractor objects [30, 26, 25], and to entirely novel objects and tasks [19, 21]. However, to the best of our knowledge, there have been no studies examining enhanced generalization to out-of-distribution object positions or end-effector poses, which is the main focus of this study. In other words, our proposed GazeBot is the first method in the general deep imitation learning framework to demonstrate the high reusability of imitated skills even under unseen object positions and end-effector poses. Although several methods exhibit similarities with GazeBot, they have not successfully demonstrated such high reusability, primarily owing to issues in the design of the action policy. Hydra [1] and SPHINX [22], for instance, segment actions into a reaching phase and a dexterous action phase, using sparse action representations similar to a bottleneck pose for the reaching motion. However, in contrast to GazeBot, these sparse representations are directly estimated by a neural network from entire images and the end-effector poses, and dexterous action prediction also relies on absolute information such as the end-effector poses or entire images. As we will see in Section IV, these design changes hinder the accurate extrapolation of the reaching motions and reduce the reusability of the learned dexterous actions when faced with unseen object positions or end-effector poses.
@@ III Gaze-based Bottleneck-aware Robot Manipulation
In this section, we propose a novel policy model named GazeBot that improves the reusability of learned skills while retaining dexterity and reactivity. We begin by explaining how gaze and bottlenecks contribute to the enhanced reusability of imitated motions (Section III-A). We then detail the data-driven action segmentation based on gaze and bottlenecks for decomposing demonstrations (Section III-B). Finally, we describe the design of the policy model (Section III-C).
@@ III-A2 Bottleneck-aware Action Segmentation
In human object manipulation, gaze position and hand movement are strongly coupled both temporally and spatially [10, 18]. As illustrated in Figure 3, a gaze-centered point cloud makes it possible to segment the movement into two distinct phases: (1) a reaching motion toward the vicinity of the gaze, and (2) a gaze-centered dexterous action. This segmentation is possible because gaze-centered vision captures only the object during the reaching phase, whereas it captures both the object and end-effector during the dexterous action phase. Here, we define the temporal boundary between these two actions as a bottleneck.
Figure 3: By observing object manipulation in gaze-centered point cloud, an action can be segmented into two phases at the bottleneck: (1) the reaching motion and (2) the gaze-centered dexterous action. The motion after the bottleneck is reusable irrespective of the object position and the initial end-effector pose.
While fully parametric methods, in which neural networks directly predict action trajectories, can achieve high dexterity and reactivity, it has been challenging to extrapolate learned behaviors to unseen object positions and end-effector poses. To address this, we restrict the use of fully parametric action prediction to the gaze-centered dexterous actions derived from the aforementioned bottleneck-aware action segmentation. In this context, the robot first reaches the bottleneck pose in a goal-fixed manner using Bézier curve approximation, and then switches to the fully parametric approach for the gaze-centered dexterous action. Here, if (A) the bottleneck estimation is independent of both object position and end-effector pose, and (B) the fully parametric gaze-centered dexterous action prediction relies solely on the relative spatial relationships between the object and end-effector (or the grasped object), it becomes possible to accurately perform the learned motions even under previously unseen object positions and initial end-effector poses without sacrificing dexterity or reactivity.
@@ III-B Data-driven Demonstration Segmentation
We first temporally segment the provided demonstrations in a data-driven manner based on gaze and bottlenecks before training an action policy. Specifically, we decompose the demonstrations into multiple sub-tasks (Section III-B1), and then further segment each sub-task into reaching motions and gaze-centered dexterous actions according to bottlenecks (Section III-B2). The corresponding action policy modules are trained using the sets of reaching motions and gaze-centered dexterous actions obtained through these processes. The decomposition into sub-tasks is performed so that each sub-task includes exactly one bottleneck by using the task decomposition method based on gaze transitions [23]. The proposed method for demonstration segmentation is summarized in Algorithm 1.
III-B1 Gaze-Based Sub-task Segmentation
A typical object manipulation task comprises multiple bottlenecks. For example, in a standard pick-and-place task, there is usually one bottleneck for the “pick” phase and another for the “place” phase. In such cases, before segmenting the motions using bottlenecks, it is necessary to decompose a sequence of object manipulation behaviors into multiple smaller sub-tasks (e.g., pick/place in this example).
As a robust and simple way to achieve such segmentation in a data-driven manner, previous work proposed a gaze-based task decomposition method for object manipulation [23]. This approach exploits the pattern of gaze during object manipulation, where the human teleoperator fixates on specific task-relevant gaze landmarks. By simply detecting transitions in this gaze behavior, the object manipulation task can be segmented so that each fixation period corresponds to a distinct sub-task. In this work, we first apply this method to decompose expert demonstrations into multiple sub-tasks.
III-B2 Bottleneck Determination
Once the demonstrations are segmented into sub-tasks, we need to determine the time step at which each sub-task reaches its bottleneck pose. To do so, we propose a bottleneck determination approach based on action predictivity from a gaze-centered point cloud. In each sub-task, predicting actions solely from a gaze-centered point cloud exhibits low accuracy when the end-effector or grasped object is not visible within the gaze-centered point cloud. Therefore, the transition from this low-predictivity phase to a high-predictivity phase can serve as the segmentation point, which we define as the bottleneck (Figure 4).
In this approach, we use a policy model $a_{t}=h(g_{t})$ that predicts the action solely from the gaze-centered point cloud $g_{t}$. This policy model is trained via behavior cloning [27], that is, by minimizing $\|h(g_{t})-a_{t}^{*}\|$ to replicate the actions in the demonstration, where $a_{t}^{*}$ denotes the expert action recorded in the demonstration. Once this policy $h$ has been trained, we compute the action-prediction loss $\|h(g_{t})-a_{t}^{*}\|$ at every time step of each demonstration. As illustrated in Figure 4, we split each sub-task based on the median of this action prediction loss, enabling entirely data-driven segmentation for every sub-task.
Figure 4: We can determine the bottleneck based on the action predictivity by training a model to predict actions solely from the gaze-centered point cloud.
@@ III-B2 Bottleneck Determination
[…]
@@
5:     Partition $[0,T^{(i)}]$ into $[s^{(i)}_{0},e^{(i)}_{0}],\dots,[s^{(i)}_{K},e^{(i)}_{K}]$
@@
$\theta^{*}=\arg\min_{\theta}\sum_{(o_{t},a_{t},g_{t})\in D}\|a_{t}-h_{\theta}(\text{crop}(o_{t},g_{t}))\|^{2}$
@@
13:for each sub-task $[s^{(i)}_{k},e^{(i)}_{k}]$ do
14:     Partition $[s^{(i)}_{k},e^{(i)}_{k}]$ into $[s^{(i)}_{k},b^{(i)}_{k}-1]$ and $[b^{(i)}_{k},e^{(i)}_{k}]$
@@ III-C Policy Design
Figure 5: GazeBot architecture. (left) The gaze prediction model is trained to estimate the gaze position across the entire image as a classification problem. (right) The action policy model achieves robust reaching motion by estimating the bottleneck pose and the shape of the trajectory up to the bottleneck, and uses a Transformer to predict the gaze-centered dexterous action in a full-parametric manner. Both actions and gaze transitions are predicted by the gaze-centered point cloud to improve the reusability.
At every time step, GazeBot first predicts the gaze position from the entire image using the gaze prediction model (Section III-C1). It then uses the gaze-centered point cloud at the predicted gaze position, along with the current end-effector pose, as inputs to the action policy model (Section III-C2 and III-C3), which reactively predicts the subsequent end-effector poses. The action policy model consists of two main modules: (i) a module that predicts gaze-centered dexterous actions solely from gaze-centered point clouds, and (ii) a module that predicts bottleneck poses based on gaze-centered point clouds and 3D gaze positions, as well as control points of first-order Bézier curves to connect the current end-effector and the predicted bottleneck pose. These modules are trained using, respectively, (i) post-bottleneck motions and (ii) reaching trajectories leading to the bottleneck pose, both obtained from demonstrations as described in the previous section III-B2.

### v3

# Enhancing Reusability of Learned Skills for Robot Manipulation via Gaze Information and Motion Bottlenecks

@@ Abstract
Autonomous agents capable of diverse object manipulations should be able to acquire a wide range of manipulation skills with high reusability. Although advances in deep learning have made it increasingly feasible to replicate the dexterity of human teleoperation in robots, generalizing these acquired skills to previously unseen scenarios remains a significant challenge. In this study, we propose a novel algorithm, Gaze-based Bottleneck-aware Robot Manipulation (GazeBot), which enables high reusability of learned motions without sacrificing dexterity or reactivity. By leveraging gaze information and motion bottlenecks—both crucial features for object manipulation—GazeBot achieves high success rates compared with state-of-the-art imitation learning methods, particularly when the object positions and end-effector poses differ from those in the provided demonstrations. Furthermore, the training process of GazeBot is entirely data-driven once a demonstration dataset with gaze data is provided. Videos and code are available at https://crumbyrobotics.github.io/gazebot.
@@ I Introduction
Recent advancements utilizing powerful neural networks such as Transformers have made deep imitation learning increasingly capable of reproducing dexterity to a certain extent [1]. However, significant issues persist regarding their generalization capabilities. Although generalization in object manipulation occurs at multiple levels, even the most fundamental aspects, such as changes in object position and the end-effector pose, are known to cause drastic reductions in success rates with variations of just a few centimeters [2].
@@
The purpose of this study is to achieve imitation that can accurately perform demonstrated object manipulations under various object positions and initial end-effector poses. Owing to the limited generalization capability of conventional imitation learning, exhaustive demonstration collection is currently required to ensure that the robot behaves correctly under various object positions and end-effector poses [3, 4]. To address this issue, we propose Gaze-based Bottleneck-aware Robot Manipulation (GazeBot), an object manipulation imitation method that enables the reuse of acquired skills even with object positions and end-effector poses not included in the provided demonstrations. As illustrated in Figure 1, the object positions and initial end-effector poses used in the demonstrations are restricted to a designated region to evaluate the reusability of the learned skills. In this setting, GazeBot can accurately perform the task not only under in-distribution (ID) conditions (i.e., within the designated region) but also under out-of-distribution (OOD) conditions (i.e., outside the designated region).
@@
To reuse skills learned within the ID domain in OOD situations, it is necessary to (1) establish an object representation that is robust to changes in object position and (2) develop an action policy architecture capable of accurate control under previously unseen end-effector poses. To achieve this, we draw inspiration from human object manipulation, where gaze on the target object not only provides a visual representation independent of the object’s absolute position but also exhibits strong gaze–hand coordination during end-effector movements such as reaching [5, 6]. Concretely, we first represent the entire field of view using a 3D point cloud and then crop a cubic region around the gaze position—referred to as gaze-centered point cloud—from the entire point cloud. By using this gaze-centered point cloud as input to the action prediction, we realize an object representation that is robust to variations in object position. Next, based on the action predictivity in the gaze-centered point cloud, we perform a data-driven action segmentation of the overall manipulation into (1) a reaching motion to the vicinity of the gaze position and (2) a gaze-centered action. We define the temporal boundary between these motions as a bottleneck, and predict the bottleneck pose, the end-effector pose at this bottleneck, from the 3D gaze position and gaze-centered point cloud. This approach allows accurate prediction of the bottleneck pose even for unseen object positions, and this prediction is independent of the current end-effector pose. Consequently, GazeBot can handle various object positions and end-effector poses by first executing a rough reaching motion to the bottleneck pose, then performing a gaze-centered action using only the gaze-centered point cloud. This approach enables the reuse of the learned motion in unseen conditions. Here, reaching the bottleneck is achieved by generating an end-effector trajectory that smoothly connects the current end-effector pose and the bottleneck pose using a first-order Bézier curve, which provides sufficient expressive power for the reaching motion while avoiding unnecessary complexity that could compromise the model’s generalization. Furthermore, GazeBot updates all actions at every step, and a fully parametric method based on a Transformer is used to directly output subsequent end-effector poses for the gaze-centered actions. This approach ensures that dexterity and reactivity are not sacrificed in the pursuit of reusability.
@@ II Related Work
Gaze-based Object Manipulation. Some works have previously proposed imitation learning methods using gaze data collected from a remote human operator during teleoperated demonstrations for action prediction [7, 8]. These gaze-based methods have exhibited advantages such as enhanced robustness by disregarding task-irrelevant objects [7] and improved dexterity by focusing on task-relevant regions of visual inputs [8]. However, because these methods rely on image cropping for gaze-centered images, they are susceptible to visual variations caused by changes in object position. In this study, we employ a gaze-centered point cloud, which provides robustness to positional changes while retaining the conventional benefits of gaze.
@@
Data-driven Action Segmentation. Segmenting actions into reaching phase and interaction phase has been proposed to improve dexterity [9], increase success rates for long-horizon tasks [10, 11], and enable high generalization capabilities [12]. Several data-driven segmentation methods have been proposed based on end-effector velocity [8] or the visibility of the end-effector within a gaze-centered image [9]. However, these approaches often fail in tasks where high dexterity is not required or when the end-effector is not visible during the interaction phase, such as manipulating with a long stick. In contrast, our approach segments motions at bottlenecks, which are determined based on action predictivity in the gaze-centered point cloud. This action predictivity-based approach offers a more general data-driven segmentation scheme compared with these previous methods.
@@
Skill Reusability. The reusability of learned skills is required across diverse levels and factors, including adaptation to a variety of object poses and unseen objects in the same category [13, 14], to changes in the environment such as varying backgrounds, camera positions, or distractor objects [15, 16], and to entirely novel objects and tasks [17]. However, to the best of our knowledge, no studies have investigated improved reusability for unseen object positions and end-effector poses without relying on object- or task-specific assumptions that compromise generality, and without sacrificing dexterity or reactivity. Our proposed GazeBot is the first deep imitation learning method to demonstrate such reusability under these challenging conditions. Although several methods exhibit similarities with GazeBot, they have not successfully demonstrated such high reusability, primarily owing to issues in the design of the action policy. Hydra [10] and SPHINX [11], for instance, segment actions into a reaching motion and a dense action, using sparse action representations similar to a bottleneck pose for the reaching motion. However, in contrast to GazeBot, these sparse representations are directly estimated by a neural network from entire images and the end-effector poses, and the dense action prediction also relies on absolute information such as the end-effector poses or entire images. As we will see in Section IV, these design changes hinder the accurate extrapolation of the reaching motions and reduce the reusability of the learned dense actions when faced with unseen object positions or end-effector poses.
@@ III-A2 Bottleneck-aware Action Segmentation
In human object manipulation, gaze position and hand movement are strongly coupled both temporally and spatially [5, 18]. As illustrated in Figure 3, a gaze-centered point cloud makes it possible to segment the movement into two distinct phases: (1) a reaching motion toward the vicinity of the gaze, and (2) a gaze-centered action. This segmentation is possible because gaze-centered vision captures only the object during the reaching phase, whereas it captures both the object and end-effector (or grasped object) during the gaze-centered action phase. Here, we define the temporal boundary between these two actions as a bottleneck.
Figure 3: By observing object manipulation in gaze-centered point cloud, an action can be segmented into two phases at the bottleneck: (1) the reaching motion and (2) the gaze-centered action. The motion after the bottleneck is reusable irrespective of the absolute object position and the initial end-effector pose.
While fully parametric methods, in which neural networks directly predict action trajectories, can achieve high dexterity and reactivity, it has been challenging to extrapolate learned behaviors to unseen object positions and end-effector poses. To address this, we restrict the use of fully parametric action prediction to the gaze-centered actions derived from the aforementioned bottleneck-aware action segmentation. In this context, the robot first reaches the bottleneck pose in a goal-fixed manner using Bézier curve approximation, and then switches to the fully parametric approach for the gaze-centered action. Here, if (A) the bottleneck estimation is independent of both object position and end-effector pose, and (B) the fully parametric gaze-centered action prediction relies solely on the relative spatial relationships between the object and end-effector (or grasped object), it becomes possible to accurately perform the learned motions even under previously unseen object positions and initial end-effector poses without sacrificing dexterity or reactivity.
@@ III-B Data-driven Demonstration Segmentation
Before training an action policy, we first temporally segment the provided demonstrations in a data-driven manner based on gaze and bottlenecks. Specifically, we decompose the demonstrations into multiple sub-tasks and then further segment each sub-task into reaching motions and gaze-centered actions according to bottlenecks. The proposed method for demonstration segmentation is summarized in Algorithm 1.
@@
$s^{(i)}_{k}\text{ (start)},\quad e^{(i)}_{k}\text{ (end)},\quad b^{(i)}_{k}\text{ (bottleneck)}$
@@
5:  Partition $0,T^{(i)}]$ into $s^{(i)}_{0},e^{(i)}_{0}],\dots,[s^{(i)}_{K},e^{(i)}_{K}$]
@@
$\theta^{*}=\arg\min_{\theta}\sum_{(o_{t},a_{t},g_{t})\sim D}\|a_{t}-h_{\theta}(\text{crop}(o_{t},g_{t}))\|^{2}$
@@
13:for each sub-task $s^{(i)}_{k},e^{(i)}_{k}$] do
14:  Partition $s^{(i)}_{k},e^{(i)}_{k}$] into $s^{(i)}_{k},b^{(i)}_{k}-1]$ and $b^{(i)}_{k},e^{(i)}_{k}$]
@@
III-B1 Gaze-Based Sub-task Segmentation
A typical object manipulation task comprises multiple bottlenecks. For example, in a standard pick-and-place task, there is usually one bottleneck for the “pick” phase and another for the “place” phase. In such cases, before segmenting the motions using bottlenecks, it is necessary to decompose a sequence of object manipulation behaviors into multiple smaller sub-tasks (e.g., pick/place in this example) so that each sub-task includes exactly one bottleneck.
As a robust and simple way to achieve such segmentation in a data-driven manner, previous work proposed a gaze-based task decomposition method for object manipulation [19]. This approach exploits the pattern of gaze during object manipulation, where the human teleoperator fixates on specific task-relevant gaze landmarks. By simply detecting transitions in this gaze behavior, the task can be segmented so that each fixation period corresponds to a distinct sub-task. In this work, we first apply this method to decompose expert demonstrations into multiple sub-tasks.
III-B2 Bottleneck Determination
Once the demonstrations are segmented into sub-tasks, we need to determine the time step at which each sub-task reaches its bottleneck pose. To do so, we propose a bottleneck determination approach based on action predictivity from a gaze-centered point cloud. In each sub-task, predicting actions solely from a gaze-centered point cloud exhibits low accuracy when the end-effector or grasped object is not visible within the gaze-centered point cloud. Therefore, the transition from this low-predictivity phase to a high-predictivity phase can serve as the segmentation point, which we define as the bottleneck.
In this approach, we use a policy model $a_{t}=h(g_{t})$ that predicts the action solely from the gaze-centered point cloud $g_{t}$. This policy model is trained via behavior cloning [20], that is, by minimizing $\|h(g_{t})-a_{t}^{*}\|$ to replicate the actions in the demonstration, where $a_{t}^{*}$ denotes the expert action recorded in the demonstration. Once this policy $h$ has been trained, we compute the action-prediction loss $\|h(g_{t})-a_{t}^{*}\|$ at every time step of each demonstration. Finally, we split each sub-task based on the median of this action prediction loss, enabling entirely data-driven segmentation for every sub-task.
@@ III-C Policy Design
Figure 4: GazeBot architecture. (left) The gaze prediction model is trained to estimate the gaze position across the entire image as a classification problem. (right) The action policy model achieves robust reaching motion by estimating the bottleneck pose and the shape of the trajectory up to the bottleneck, and uses a Transformer to predict the gaze-centered action in a full-parametric manner. Both actions and gaze transitions are predicted by the gaze-centered point cloud to improve the reusability.
At every time step, GazeBot first predicts the gaze position from the entire image using the gaze prediction model. It then uses the gaze-centered point cloud at the predicted gaze position, along with the current end-effector pose, as inputs to the action policy model, which reactively predicts the subsequent end-effector poses.